OpenStudy (anonymous):

If a+b=111 and a-b=7 then ab=?

OpenStudy (anonymous):

a+b=11

OpenStudy (anonymous):

a+b=11, a-b=7 (a+b)+(a-b)=11+7 2a=18 a=9, then b=2 ab=9*2=18